# Bridging the Long-Term Gap: A Memory-Active Policy for Multi-Session Task-Oriented Dialogue

### v1

@@ 5.1 Experimental Setups
Evaluation Metrics. We use four core metrics to evaluate model performance: GPT-4 score, Joint Goal Accuracy (JGA), Dialogue Turn Efficiency (DTE), and Success Rate (S.R.). GPT-4 Score (1–10)  [^3] GPT4-as-the-judge prompts can be found in Appendix A.4 reflects response quality based on fluency, coherence, and informativeness. JGA measures slot prediction accuracy, DTE captures the number of turns required to complete a task, and S.R. indicates whether the user goal is achieved. To support analysis, we report auxiliary metrics including Recall@k for memory retrieval accuracy, Slot Accuracy for value correctness, and BLEU and ROUGE for generation overlap. Human evaluation further assesses Accuracy, Informativeness, and Coherence, with A.I.C. denoting their average. Annotation details are provided in Appendix C.3.
Model Setting GPT4 JGA DTE S.R.
LLaMA3-8B w/o MAP 4.89 0.64 5.37 0.82
w/ MAP 6.39 0.63 3.46 0.92
[…]
w/ MAP 6.81 0.66 4.31 0.87
[…]
w/ MAP 6.48 0.80 1.21 1.00
[…]
w/ MAP 7.14 0.70 3.19 0.99
Table 3: Performance comparison of general-purpose LLM models with and without long-term memory integration. The w/o MAP setting uses full-context prompting, feeding the entire dialogue history as input, while w/ MAP leverages memory active policy to retrieve and utilize relevant long-term memory.
Baselines. We comprehensively evaluate MAP against three baseline categories:
  • General-purpose LLMs. We evaluate prompting-based dialogue performance with instruct models including LLaMA3-8B Touvron et al. (2024), Qwen2.5-7B Team (2024c), Mistral-7B Team (2024a), and GPT-4o-mini Team (2024b).
  • Traditional Task-Oriented Dialogue Systems. To assess MAP in structured DST scenarios, we compare with task-specific fine-tuned BERT-DST Chao and Lane (2019), AutoTOD Xu et al. (2024a), and LDST Feng et al. (2024). Among these, AutoTOD incorporates an external memory module to track user goals across turns.
  • Long-term Summarization Baseline. We additionally implement a summarization-based baseline inspired by ChatCite Li et al. (2024c), which produces concise history summaries for each session. The same inference model (GPT-4o-mini) is used across all settings to ensure fairness.
To ensure a fair assessment of generalization, all models are evaluated on a held-out multi-session test set that is excluded from all training processes.
5.2 Main Results
Comparision with General-purpose LLMs. We conduct the experiments comparing full context prompting and our MAP framework in the metric of GPT4, JGA, DTE, and S.R. As shown in Table 3, MAP demonstrates consistent performance gains over baseline prompting methods. For instance, applying MAP to Mistral-7B increases JGA from 0.73 to 0.80 and S.R. from 0.83 to 0.87. Notably, LLaMA3-8B, Qwen-7B, and GPT-4o-mini also show significant improvements in both JGA and S.R. when integrated with MAP. In terms of response quality, GPT-4 score rises notably for all models; for example, LLaMA3-8B achieves the largest gain, from 4.89 to 6.39. Regarding DTE, MAP considerably shortens the required turns, with reductions of 35.6% for LLaMA3-8B, 12.6% for Qwen-7B, 52.0% for Mistral-7B, and 47.1% for GPT-4o-mini. These results demonstrate that integrating long-term memory enhances both response quality and conversation efficiency.
Model GPT4 JGA DTE S.R.
Bert-DST - 0.067 - -
LDST - 0.234 - -
AutoTOD 6.49 0.440 7.80 0.81
@@ 5.2 Main Results
MAP 7.14 0.698 3.19 0.99
Table 4: Performance comparison of traditional TOD models, summary-based methods, and MAP. Models marked with focus on DST only. indicates simplified AutoTOD pipeline. The ChatCite represents a long-term memory baseline using dialogue summarization.
Model Confirmation Multi-Turn
w/o MAP w/ MAP w/o MAP w/ MAP
LLaMA3-8B 1.64 1.99 1.60 2.03
Qwen-7B 1.46 1.88 1.48 1.77
Mistral-7B 1.79 1.99 2.04 2.18
GPT4o-mini 1.86 2.27 1.72 1.85
Table 5: Human evaluation results based on the average A.I.C., which is the mean of Accuracy, Informativeness, and Coherence. w/ denotes with, w/o denotes without.
Comparison with Traditional TOD and Long-term Summarization Models. As no prior model explicitly targets multi-session TOD, we compare MAP with traditional DST baselines (BERT-DST, LDST) and full-pipeline systems (AutoTOD). As shown in Table 4, MAP substantially outperforms these models, achieving a JGA of 0.698 (vs. 0.440 for AutoTOD) and reducing DTE from 7.8 to 3.19. We also evaluate against a summarization-based approach (ChitChat), where MAP again achieves superior results: higher GPT-4 score (7.14 vs. 6.98), better JGA (0.70 vs. 0.66), shorter DTE (3.19 vs. 4.71), and improved success rate (0.99 vs. 0.84). These results highlight the effectiveness of MAP’s memory integration and proactive strategy in enhancing both accuracy and interaction efficiency.
Human Evaluation. We conduct human evaluation to further assess the effectiveness of the MAP structure across two key dialogue settings: (1) confirmation-type responses after memory-guided dialogue planning, and (2) multi-turn dialogues under the proactive response strategy. Evaluators are instructed to rate responses based on Accuracy (binary), Informativeness (Likert 0–3), and Coherence (0–3). A weighted average score (A.I.C.) is used to summarize overall performance. As shown in Table 5, MAP consistently improves perceive response quality across all metrics. All evaluations are conducted in a blind review setup. Further annotation guidelines, examples, and scoring distributions are provided in Appendix C.3.
5.3 Extended Evaluation
To examine the applicability of our QA memory framework beyond multi-session TOD, we evaluate MAP on two widely-used DST benchmarks, SGD and MultiWOZ2.2. Although both benchmarks target dialogue state tracking (DST), they differ in annotation formats and domain complexity, leading to different baseline sets (Table 7). On SGD, MAP achieves a state-of-the-art JGA of 0.846, surpassing strong baselines such as LDST Feng et al. (2023), GOLOMB Gulyaev et al. (2020), SGP-DST Ruan et al. (2020), and TS-DST Du et al. (2022), and performs comparably to LDST on AGA. On MultiWOZ2.2, MAP* attains a JGA of 0.879, significantly outperforming prior models including TRADE Wu et al. (2019a), TripPy Heck et al. (2020), and SDP-DST Lee et al. (2021). We attribute the superior performance to QA memory’s ability to capture slot dependencies more effectively in smaller domain settings, confirming its adaptability and robustness across datasets.
5.4 Ablation Study
We conduct ablations to evaluate contributions of key MAP components, including the judger-refinement module and proactive response strategy.
Figure 4: Comparison of Hybrid RAG (blue) vs. MAP (purple) across four LLMs on two metrics: JGA and DTE.
Judgement-Refinement. We conduct ablation experiments to analyze the contribution of the judger and refinement components in MAP. Compared with a strong retrieval-based baseline, we evaluate MAP with and without the judger-refinement module, using the same Hybrid RAG-based memory activation. As shown in Figure 4, MAP with judger-refinement consistently outperforms the baseline across model backbones. On Qwen2.5-7B, JGA rises from 0.41 to 0.74, and on GPT-4o-mini, DTE drops from 4.30 to 3.19, highlighting the benefit of goal-aware filtering and structured memory usage.
To isolate the effect of the Memory Judger, we remove it from the pipeline while keeping the refinement step. The judger uses Chain-of-Thought (CoT) prompting to infer missing task slots and assess the alignment of QA pairs with current dialogue goals. This reasoning-guided filtering improves memory relevance, yielding a Recall@5 gain of 7.7% on average across models (Appendix Figure 6); for instance, text-embedding-3-small improves from 0.792 to 0.832.
To assess the effect of the Refinement module, we remove it while retaining the judger. Refinement restructures selected QA pairs by discarding auxiliary questions and preserving only core answers aligned with intent. As shown in Table 6, removing Refinement causes performance to drop: JGA decreases from 0.70 to 0.64, GPT-4 Score from 7.14 to 6.94, and Success Rate from 0.99 to 0.88, while DTE increases from 3.19 to 4.33. These results demonstrate that both components are essential for precise memory integration and efficient multi-session dialogue.
Proactive Response Strategy. MAP identifies incomplete or inconsistent slot values by comparing predicted responses with task goals, and proactively engages the user to resolve them. The effectiveness of this strategy is reflected in the observed improvement from 0.68 slot accuracy prior to correction to a task success rate of 0.99 after correction. These results demonstrate its critical role in early error resolution and goal fulfillment. Detailed statistics are provided in Appendix C.2.
Model Variant GPT-4 Score JGA DTE S.R.
MAP 7.14 0.70 3.19 0.99
- w/o refinement 6.94 0.64 4.33 0.88
Table 6: Impact of memory refinement in the memory activation module. Refinement improves JGA, GPT-4 score, and overall task success.
@@ 5.4 Ablation Study
Table 7: Results of different methods on SGD and MultiWOZ 2.2. $\mathrm{MAP}^{*}$ is a single-session variant of MAP, where the Memory Judger is disabled while retaining the QA memory.
5.5 Case Study
In our case study, we compared four methods for generating confirmation responses: (1) Direct Prompting with full conversation history, (2) Hybrid RAG retrieving relevant dialogue history, (3) Hybrid RAG retrieving a summary of the conversation, and (4) MAP with intention-based QA memory. As shown in Figure 5, Direct Prompting introduces noise, while Hybrid RAG variants often omit key task slots due to retrieval inaccuracies. In contrast, MAP consistently produces accurate, concise responses with fewer errors. These findings indicate that MAP with intention-based QA memory better preserves key task information and reduces errors in confirmation generation than other baselines. More details are provided in Appendix D.
6 Related Works
6.1 Task-Oriented Dialogue Dataset
TOD datasets are typically constructed via either Machine-to-Machine (M2M) Shah et al. (2018); Rastogi et al. (2020) or Wizard-of-Oz (WOz) setups Wen et al. (2017); Budzianowski et al. (2018). M2M datasets (e.g., SGD, STAR) provide schema-driven task flows, while WOz-based datasets (e.g., MultiWOZ, FRAMES) offer more natural but annotation-heavy dialogues. Recent efforts aim to improve realism and domain diversity Zhang et al. (2022); Hu et al. (2023); Dai et al. (2022); Xu et al. (2024b); Li et al. (2024b), yet existing benchmarks primarily assume single-session tasks. There remains a notable gap in datasets designed for multi-session TOD, where tracking long-range goals and user intents is essential.
@@ 6.1 Task-Oriented Dialogue Dataset
6.2 Task-Oriented Dialogue Systems
Traditional TOD systems adopt modular pipelines for NLU, DST, and response generation Wu et al. (2019b); Peng et al. (2018), later unified into end-to-end models trained on annotated dialogues Wen et al. (2017); Wang et al. (2020). With the rise of LLMs, recent work explores their use in zero-shot and fine-tuned TOD Madotto et al. (2021); Bang et al. (2023), often achieving strong results on intent recognition and slot filling. In parallel, long-term memory (LTM) methods such as ChitChat Li et al. (2024c), MemoryBank (Zhong et al., 2024), and LoCoMo Maharana et al. (2024) support extended context retention through summarization or heuristic filtering, but lack structured memory aligned with task goals. Most assume single-session dialogues and overlook challenges in maintaining multi-session goal continuity. This work addresses these gaps by introducing a memory-active policy for long-range, goal-aware tracking.
7 Conclusion
This study introduces a multi-session TOD task and the MS-TOD dataset, which features diverse task goals and structured memory banks across sessions. To address the challenges of long-term context integration, we propose MAP, a memory-augmented framework that combines dialogue planning with a proactive response strategy for efficient task completion. Experimental results show that MAP significantly reduces dialogue turns, improves response quality, and boosts task success rate, outperforming both direct prompting and existing long-term retrieval methods.
Limitation
While our model demonstrates effectiveness on the current dataset, several limitations remain. First, our experiments are limited to locally deployable LLM models, and we have not explored the potential benefits of scaling to larger models, which may yield further improvements. Second, our approach does not incorporate external knowledge bases or internet search functionality, which could enhance contextual understanding and factual accuracy. Lastly, the model’s generalizability to broader domains and more complex real-world scenarios remains untested, necessitating further evaluation across diverse datasets and tasks. Future work will address these limitations by expanding model scalability, integrating external knowledge sources, and conducting more comprehensive evaluations.
Ethics Statement
Our research enhances multi-session task-oriented dialogue through memory-augmented processing while complying with ethical standards. All datasets are public and contain no personally identifiable information; no user-sensitive data are collected, and no human subjects are involved. Model evaluations are conducted by three trained research assistants, each paid \$20/hour, above the local average. We acknowledge risks including misinformation, algorithmic bias, and issues specific to long-term memory, such as retaining outdated user preferences or privacy concerns from cross-session data accumulation. These are addressed through evaluation safeguards, memory control mechanisms, and design principles promoting transparency, fairness, and accountability.
Appendix A Prompts
A.1 Prompt of dialogue generation
@@
We designed a multi-session dialogue prompt (as shown in Figure 7) that generates multi-session dialogue data based on input dialogue intent, task goal, and target session count. Additionally, during the generation process, we annotate whether each utterance is a confirmation response. These annotations, after manual verification, will be used in the main experiment for confirmation-type response generation.
A.2 Prompt of Task Slot Query Generation
@@
During the evaluation process, we design a prompt (as shown in Figure 8) that generates a query corresponding to the missing task attributes based on the current dialogue context and task objectives. The input to this prompt is the dialogue context history and the generated task objectives. This query is then used as input to the memory judger to assist in selecting QA memory units that align with the task objectives.
A.3 Prompts of Confirmation Response Generation
@@
In the evaluation process, we employed a confirmation-type response generation approach to assess the integration performance of multi-session memory in task-oriented dialogues. We designed the prompt as shown in Figure 9, which leverages the dialogue context, task objectives, and activated memory units to generate responses.
A.4 Prompts of GPT4 Evaluation
@@
During the evaluation process, we employed a GPT-4 prompt (as shown in Figure 11) to assess the quality of confirmation-type responses. This prompt evaluates the response holistically from four perspectives: requirement alignment, content accuracy, language quality, and comparison to the reference answer. The input to this prompt includes the dialogue history, task objectives, the reference response, and the model-generated response. This design ensures that the evaluation of the response is not solely based on the dataset’s reference reply but also takes into account multiple factors such as whether the task objectives are met and the overall quality of the response. Such an evaluation approach is more comprehensive.
@@
Table 8: MS-TOD Subset Statistics for Memory Judger Training.
A.5 Prompts of Dialogue State Tracking
we used a prompt modified from Heck et al. (2023b) (as shown in Figure 10) that generates the dialogue state for each user turn in the dialogue. Let
@@ A.5 Prompts of Dialogue State Tracking
where P is the task description which provides the model with instructions for how to process a dialogue between a system M and a user U. In contrast to Heck et al. (2023b), P does not include the detailed description for slots to challenge ChatGPT’s ability to understand the meaning of the slots. Apart from that, ChatGPT often generated answers with excessively detailed explanations, deviating from the expected response format. To address this issue, a prompt that includes "No explanation!" as an instruction to ChatGPT not to provide detailed explanations was introduced Feng et al. (2023) and we added this to our prompt.
Dataset Task Type Multi- Session? Grounded Memory? User Intention? Retrieval Support? Memory Format
MultiWOZ (Hu et al., 2023) TOD
SGD (Rastogi et al., 2020) TOD schema
TOAD (Liu et al., 2024b) TOD
LUCID (Stacey et al., 2024) TOD latent goal
MSC (Xu et al., 2022) OD dialogue history
CC (Jang et al., 2023) OD persona/dialogue history
MemoryBank (Zhong et al., 2024) OD dialogue history
LoCoMo (Maharana et al., 2024) OD dialogue history
LongMemEval (Wu et al., 2025) OD dialogue history
@@
Appendix B Dataset
B.1 Dataset for Memory Judger
To ensure that the memory judger generalizes across different domains and scenarios, we generated the training dataset(as shown in Table 8) using the same method described in the main text. The dataset spans 16 domains, 4,534 task goals, and 13,411 dialogues, involving a total of 565 individuals, each with an average of 6.24 intentions. Beyond training the memory judger, this dataset can also serve as an alternative evaluation set for broader benchmarking.
B.2 Dataset Structure
MS-TOD encompasses multiple individual task-oriented dialogue datasets, each consisting of several sessions. We present an example of one session (as shown in Figure 12) from an individual. This session includes a $session\_id$, where a larger value indicates a more recent timestamp. The domain represents the specific field or area of the dialogue. The $reference\_dialogue\_id$ corresponds to the $dialogue\_id$ in the original SGD dataset that shares the same task objective. The $exist\_confirmation$ indicates whether the session contains a confirmation-type response and whether it is an evaluation target. The intent represents the specific purpose or goal of the dialogue. The content stores the actual dialogue text. The $task\_goal$ includes task slots and their corresponding attribute values. Each individual contains dozens of session data structured as described above.
B.3 Human Validation Protocol

### v2

@@ Introduction
We use four core automatic metrics and human evaluation to evaluate response performance: 1) GPT-4 score, (1–10)  [^3] GPT-4-as-the-judge prompts can be found in Appendix evaluates response quality in terms of fluency, coherence, and informativeness; 2) Joint Goal Accuracy (JGA) measures slot prediction accuracy; 3) Dialogue Turn Efficiency (DTE) captures the number of turns required to complete a task, and 4) Success Rate (S.R.) indicates whether the user goal is achieved. 5) Human evaluation further assesses Accuracy, Informativeness, and Coherence, with A.I.C. denoting their average.
Model Setting GPT-4 JGA DTE S.R. A. I. C. A.I.C.
LLaMA3-8B FCP 4.89 0.64 5.37 0.82 0.56 1.47 1.74 1.26
MemGuide 6.39 0.63 3.46 0.92 0.61 1.98 2.16 1.58
Qwen-7B FCP 6.26 0.66 4.93 0.83 0.43 1.24 1.85 1.17
MemGuide 6.81 0.66 4.31 0.87 0.54 1.70 2.30 1.51
Mistral-7B FCP 6.20 0.73 2.52 1.00 0.58 1.63 1.99 1.40
MemGuide 6.48 0.80 1.21 1.00 0.61 2.06 2.08 1.58
GPT-4o-mini FCP 6.93 0.67 6.03 0.88 0.62 1.83 1.90 1.78
MemGuide 7.14 0.70 3.19 0.99 0.65 2.38 2.48 2.17
Table 3: Combined results comparing FCP and MemGuide across LLMs. GPT-4, JGA, DTE, and S.R. are automatic metrics; A., I., C., and A.I.C. are human evaluation metrics for accuracy, informativeness, coherence, and their average score.
Baselines.
We evaluate MemGuide against three representative categories of baselines: 1) General-purpose LLMs. We assess full-context prompt (FCP)-based dialogue performance using instruction-tuned models including LLaMA3-8B (Touvron et al. 2024), Qwen2.5-7B (Team 2024c), Mistral-7B (Team 2024a), and GPT-4o-mini (Team 2024b). 2)Traditional Task-Oriented Dialogue Systems. To evaluate MemGuide under structured DST conditions, we include task-specific baselines such as BERT-DST (Chao and Lane 2019), LDST (Feng et al. 2024), and AutoTOD (Xu et al. 2024a), where the latter incorporates an external memory module for cross-turn goal tracking. While these models were not originally designed for multi-session scenarios, they represent the strongest available TOD pipelines when adapted to this setting. 3) Long-term Summarization. We implement a summarization-based baseline inspired by ChatCite (Li et al. 2024c), which condenses session histories into concise summaries used during inference.
Main Results
Comparision with General-purpose LLMs. Compared to FCP, MemGuide leverages the same underlying LLM as a memory reader to retrieve and utilize relevant long-term memory, yielding substantial improvements across critical metrics. As illustrated in Table 3, our approach consistently outperforms FCP across all tested models, demonstrating robust gains in task accuracy, response quality, and interaction efficiency. For example, when using Mistral-7B as the LLM Reader, MemGuide increases JGA from 0.73 to 0.80 and reduces DTE from 2.52 to 1.21, representing a 52% reduction. LLaMA3-8B achieves the largest improvement in GPT-4 score (from 4.89 to 6.39), while GPT-4o-mini reduces dialogue turns from 6.03 to 3.19, corresponding to a 47.1% decrease.. Similar gains are observed with Qwen-7B and other models. These consistent gains across models confirm that integrating memory-guided reasoning with the same base model enhances not only task accuracy but also response relevance and interaction fluency, validating the effectiveness of intent-aligned memory selection in multi-session task-oriented dialogue.
Model GPT-4 JGA DTE S.R.
Bert-DST - 0.07 - -
LDST - 0.23 - -
AutoTOD 6.49 0.44 7.80 0.81
@@
MemGuide 7.14 0.70 3.19 0.99
Table 4: Results of traditional TOD models, summary-based methods, and MemGuide. Models marked with focus on DST only. indicates a simplified AutoTOD pipeline.
@@
Table 5: Results of different methods on SGD and MultiWOZ $2.$ $\mathrm{MemGuide}^{*}$ is a single-session variant of MemGuide, where the missing slot guided filtering is disabled while retaining the QA memory.
Comparison with Traditional TOD and Summarization Baselines. As existing models are not explicitly designed for multi-session TOD, we compare MemGuide with two representative categories: (1) traditional Dialogue State Tracking (DST)-focused models (BERT-DST, LDST, AutoTOD), and (2) a summarization-based approach inspired by ChatCite. As shown in Table 4, MemGuide consistently outperforms all baselines across key metrics. Compared to AutoTOD, it increases JGA from 0.440 to 0.698 and reduces DTE from 7.80 to 3.19. Against the summarization baseline, MemGuide achieves clear improvements across all metrics: the GPT-4 score increases from 6.59 to 7.14, JGA improves from 0.66 to 0.70, DTE decreases from 4.71 to 3.19, and the success rate rises from 0.84 to 0.99. These results demonstrate the effectiveness of MemGuide in improving both task performance and dialogue efficiency.
Human Evaluation. We conduct human evaluation to assess the effectiveness of MemGuide in generating confirmation-type responses after memory-guided dialogue planning. Human annotators rate each response along three dimensions: Accuracy (binary), Informativeness (scored from 0 to 3), and Coherence (scored from 0 to 3). The average of these scores, denoted as A.I.C., provides an overall measure of perceived response quality. As shown in Table 4, MemGuide consistently improves human-judged quality across all metrics. All evaluations are conducted under a blind review protocol.
Generalization to Single-Session DST Tasks. To assess the generalization of MemGuide in single-session settings, we focus on dialogue state tracking (DST), a core task that supports downstream components like policy planning and response generation. DST is a widely used and well-defined task that depends on context understanding and supports downstream dialogue components. We evaluate it on two widely-used single-session DST benchmarks, SGD and MultiWOZ2.2. While both benchmarks focus on DST, differences in annotation schemes and domain coverage result in distinct baseline configurations across SGD and MultiWOZ2.2 (Table 5). On SGD, MemGuide achieves a state-of-the-art JGA of 0.846, surpassing strong baselines such as LDST (Feng et al. 2023), GOLOMB (Gulyaev et al. 2020), SGP-DST (Ruan et al. 2020), and TS-DST (Du et al. 2022), and performs comparably to LDST on Average Goal Accuracy (AGA) (Rastogi et al. 2020). On MultiWOZ2.2, MemGuide* attains a JGA of 0.879, significantly outperforming prior models including TRADE (Wu et al. 2019b), TripPy (Heck et al. 2020), and SDP-DST (Lee, Cheng, and Ostendorf 2021). We attribute the superior performance to QA memory’s ability to capture slot dependencies more effectively in smaller domain settings, confirming its adaptability and robustness across datasets.
Setting w/ Raw History w/ Intent-QA Memory
LLaMA3-8B 5.09 6.34
Qwen-7B 6.38 6.56
Mistral-7B 5.86 6.71
GPT-4o-mini 7.01 7.14
Table 6: Comparison of GPT-4 scores using retrieved history vs. intent-QA memory across different LLM settings.
Ablation Study
Effect of intent-aligned Retrieval. Table 6 shows that our intent-aligned retrieval with structured QA memory consistently improves response quality, outperforming unstructured baselines by up to 1.29 points (e.g., LLaMa3-8B: 5.05 → 6.34). We attribute this enhancement to the alignment between the retrieved structured context and the model’s generative reasoning process. By providing intent-anchored QA pairs, the system can reason over task-relevant content with greater precision and mitigated ambiguity, thereby corroborating the theoretical premise that structured supervision is key to enhancing long-context utility in dialogue systems.
Effect of Missing-Slot Guided Filtering. We assess the impact of the missing-slot guided filtering module by removing it while retaining the same hybrid RAG retrieval. As shown in Figure 4, the absence of filtering results in a significant performance drop: JGA on Qwen2.5-7B drops from 0.74 to 0.41, and DTE on GPT-4o-mini increases from 3.19 to 4.30. This highlights the critical role of fine-grained memory selection in both accuracy and interaction efficiency. This module operates in two stages: the CoT reasoner identifies missing task slots, which guide the memory filter to prioritize QA pairs that fill these gaps. This joint reasoning and filtering strategy significantly improves retrieval quality: compared to semantic-only retrieval, Recall@5 improves by 7.7% on average, raising the performance of text-embedding-3-small from 0.792 to 0.832. By bridging retrieval with task-specific gaps, the filter prioritizes slot-value pairs that are more likely to advance the dialogue.
Figure 4: Comparison of Hybrid RAG (blue) vs. MemGuide (purple) across four LLMs on two metrics: JGA and DTE.
@@
Case Study
In our case study, we compared four methods for generating confirmation responses: (1) FCP, (2) Hybrid RAG, (3) ChatCite, and (4) MemGuide. As shown in Figure 5, FCP often introduces irrelevant or outdated content due to unfiltered long context, while Hybrid RAG and ChatCite frequently miss critical slot values such as dates or preferences, reflecting retrieval noise or lack of task-specific reasoning. In contrast, MemGuide consistently produces accurate and concise responses by combining intent-aligned retrieval with missing-slot guided filtering. Notably, it achieves higher slot coverage and fluency than other methods, as verified by both GPT-4 scores and manual annotation. These results highlight the value of intent-aware retrieval and task-specific filtering for enhancing response quality in multi-session TOD.
Conclusion
We present MemGuide, a two-stage memory-guided framework for multi-session LLM agents. By combining intent-aligned retrieval with missing-slot guided filtering, MemGuide enables task-aware, slot-specific memory selection that surpasses traditional semantic similarity. Evaluated on MS-TOD, our novel benchmark for multi-session TOD, MemGuide significantly improves task success, shortens dialogues, and enhances interaction coherence. These results confirm that structured memory supervision and goal-aware reasoning are critical for developing effective LLM agents.
Appendix A Appendix A
Prompt of dialogue generation
@@ Appendix A Appendix A
We designed a multi-session dialogue prompt (as shown in Figure A.1) that generates multi-session dialogue data based on input dialogue intent, task goal, and target session count. Additionally, during the generation process, we annotate whether each utterance is a confirmation response. These annotations, after manual verification, will be used in the main experiment for confirmation-type response generation.
Prompt of Task Slot Query Generation
@@
During the evaluation process, we design a prompt (as shown in Figure A.2) that generates a query corresponding to the missing task attributes based on the current dialogue context and task objectives. The input to this prompt is the dialogue context history and the generated task goal. This query is then used as input to the memory judger to assist in selecting QA memory units that align with the task objectives.
Prompts of Confirmation Response Generation
@@
In the evaluation process, we employed a confirmation-type response generation approach to assess the integration performance of multi-session memory in task-oriented dialogues. We designed the prompt as shown in Figure A.3, which leverages the dialogue context, task objectives, and activated memory units to generate responses.
Prompts of GPT4 Evaluation
@@
During the evaluation process, we employed a GPT-4 prompt (as shown in Figure A.5) to assess the quality of confirmation-type responses. This prompt evaluates the response holistically from four perspectives: requirement alignment, content accuracy, language quality, and comparison to the reference answer. The input to this prompt includes the dialogue history, task objectives, the reference response, and the model-generated response. This design ensures that the evaluation of the response is not solely based on the dataset’s reference reply but also takes into account multiple factors such as whether the task objectives are met and the overall quality of the response. Such an evaluation approach is more comprehensive.
@@
Table 7: MS-TOD Training Dataset Statistics for Memory Fileter.
Prompts of Dialogue State Tracking
we used a prompt modified from (Heck et al. 2023) (as shown in Figure A.4) that generates the dialogue state for each user turn in the dialogue. Let
@@
where P is the task description which provides the model with instructions for how to process a dialogue between a system M and a user U. In contrast to (Heck et al. 2023), P does not include the detailed description for slots to challenge ChatGPT’s ability to understand the meaning of the slots. Apart from that, ChatGPT often generated answers with excessively detailed explanations, deviating from the expected response format. To address this issue, a prompt that includes ”No explanation!” as an instruction to ChatGPT not to provide detailed explanations was introduced (Feng et al. 2023) and we added this to our prompt.
Dataset Task Type
Multi-
Session?
Grounded
Memory?
User
Intention?
Retrieval
Support?
Memory
Format
MultiWOZ (Hu et al. 2023) TOD
SGD (Rastogi et al. 2020) TOD schema
TOAD (Liu et al. 2024b) TOD
LUCID (Stacey et al. 2024) TOD latent goal
MSC (Xu, Szlam, and Weston 2022) OD dialogue history
CC (Jang, Boo, and Kim 2023) OD persona/dialogue history
MemoryBank (Zhong et al. 2024) OD dialogue history
LoCoMo (Maharana et al. 2024) OD dialogue history
LongMemEval (Wu et al. 2025) OD dialogue history
@@
Appendix B Appendix B
Training Dataset for Memory Filter
To ensure that the memory filter generalizes across different domains and scenarios, we generated the training dataset(as shown in Table 7) using the same method described in the main text. The dataset spans 16 domains, 4,534 task goals, and 13,411 dialogues, involving a total of 565 individuals, each with an average of 6.24 intentions. Beyond training the memory filter, this dataset can also serve as an alternative evaluation set for broader benchmarking.
Dataset Structure
MS-TOD encompasses multiple individual task-oriented dialogue datasets, each consisting of several sessions. We present an example of one session (as shown in Figure A.6) from an individual. This session includes a session_idsession\_id, where a larger value indicates a more recent timestamp. The domain represents the specific field or area of the dialogue. The reference_dialogue_idreference\_dialogue\_id corresponds to the dialogue_iddialogue\_id in the original SGD dataset that shares the same task objective. The exist_confirmationexist\_confirmation indicates whether the session contains a confirmation-type response and whether it is an evaluation target. The intent represents the specific purpose or goal of the dialogue. The content stores the actual dialogue text. The task_goaltask\_goal includes task slots and their corresponding attribute values. Each individual contains dozens of session data structured as described above.
Human Validation Protocol